Question 1 (2 points, 1 point for each correct answer) A research group discovered a new enzyme happyase that catalyzes chemical reaction SAD STUDENT HAPPY STUDENT Researchers characterized this enzyme and found that at [E,-10 nM this enzyme has Vmax at 2.5 μM/s. It was also experimentally measured that at [SAD STUDENT)-0.05 mM this enzyme has Vo -500 nM/s. What are the kat and Km of the happyase enzyme? Make sure to write down units of measure! Answers: keat- Question 2 (1 point): sample of DNA isolated from bacterium X contains 17.5% of adenine bases. What percentages of guanine, thymine, and cytosine bases would you expect to find in the same sample of DNA? Answer:
Answer #1

Answer:

Based on the given data:

  • Etot = 10nM
  • Vmax = 2.5uM/s = 2500nM/s (*1uM = 1000nM)
  • At [SAD_STUDENT] = 0.05mM, V0 =500 nM/s
  • Kcat and Km = ?
  • Using the Michaelis Menten equation:

V_0 = \frac{V_{max}*[S]}{K_m + [S]}

K_m = \frac{V_{max}*[S]- V_0[S]}{V_0}

K_m = \frac{V_{max}*[S]- V_0[S]}{V_0} = \frac{2500*0.05- 500*0.05}{500} = 0.2mM

K_{cat} = \frac{V_{max}}{E_{tot}} = 2500/10 =250 s^{-1}